OpenStudy (imnotgoodatmath):

1. Solve -4x + 15 > 59 2. 3x - 9 + x > 27 3. 4 (5m + 6) - 8 < 336

OpenStudy (anonymous):

1. x<-11

OpenStudy (anonymous):

1. x<-11 2. x>9 3.

OpenStudy (anonymous):

3. m<16

OpenStudy (anonymous):

2.x>9

OpenStudy (anonymous):

3.m<16

OpenStudy (anonymous):

\(\large \color{blue}{1.}\)\(\large -4x + 15 > 59\\\large -4x>59-15=44\\\large 4x<-44\\\large x<-11\) \(\large \color{blue}{2.}\)\(\large 3x - 9 + x > 27\\\large 4x>27+9=36\\\large x>9\) \(\large \color{blue}{3.}\)\(\large 4 (5m + 6) - 8 < 336\\\large 20m+24-8<336\\\large 20m<336-16=320\\\large m<16\)